# Centro de Física do Porto e Departamento de Física e Astronomia, Faculdade de Ciências da Universidade do Porto, Rua do Campo Alegre 687, 4169-007, Porto, Portugal τCPHT, CNRS, École Polytechnique, Institut Polytechnique de Paris, Palaiseau, France

### v1

@@ 2.2 Continuum limit and candidate fixed points
  • $c=8/7\approx 1.14$ corresponding to the $\mathbb{Z}_{5}$ parafermion and its bosonized coset $SU(2)_{5}/U(1)$ Fateev:1985mm ; Dotsenko:2002gs .
@@
  • $c=81/70\approx 1.15$ corresponding to the third $\mathcal{N}=1$ super-Virasoro minimal model and its bosonized coset $SU(2)_{2}\times SU(2)_{3}/SU(2)_{5}$ Goddard:1986ee .
[…]
Generally, one should always keep in mind discrete orbifolds of any CFT, since these have the same central charge. These are often related to different modular invariants. Based on this list, another possible candidate for the infrared fixed point is the bosonic coset
@@
while the other theories do not realize the symmetries in any obvious way. In the coset (35), one of the two invertible $\mathbb{Z}_{2}$ symmetries is manifest, as well as one of the Fibonacci lines: they are both Verlinde lines Verlinde:1988sn . It is possible that a second Fibonacci and $\mathbb{Z}_{2}$ lines exist as non-Verlinde lines, potentially coming from a $U(1)\subset SU(2)$ Wilson type line Chang:2018iay . To analyse this candidate more carefully, we look at the low lying spectrum of scalar operators. The coset primaries are labeled by triplets of $SU(2)$ representations $(\ell_{2},\ell_{3},\ell_{5})$, where $\ell_{i}\in\{0,1/2,\dots,k/2\}$, subject to the identification $\ell_{i}\equiv k_{i}/2-\ell_{i}$. Furthermore, not all triplets correspond to allowed representations, and are further subject to the constraint
$\ell_{2}+\ell_{3}-\ell_{5}\in\mathbb{Z}\,.$ (36)
As is well known, the stress-tensor of the coset model is given by the difference of the numerator and denominator theories’ stress tensor (i.e. $T_{\textrm{coset}}=T_{2}+T_{3}-T_{5}$), and therefore the chiral dimensions of coset primaries satisfy
$h_{(\ell_{2},\ell_{3},\ell_{5})}=h_{\ell_{2}}+h_{\ell_{3}}-h_{\ell_{5}}+n\,,n\in\mathbb{N}\,,$ (37)
where the addition of a positive integer might be needed since the primary does not always appear in the ground state of the denominator representation. Here, $h_{\ell_{k}}$ are the weights of primaries in the $SU(2)_{k}$ WZW model WESS197195 ; Witten:1983tw and are given by
$h_{\ell_{k}}=\frac{\ell_{k}(\ell_{k}+1)}{k+2}\,.$ (38)
Given the finite list of chiral primaries, we can try to build a modular invariant by observing if any chiral dimensions differ by an integer, in which case a non-diagonal modular invariant might exist. For our case, no such integer differences occur, [^10] Instead, one does find half-integer differences, and in particular a chiral dimension $3/2$. This is because the chiral theory admits a fermionic partition function with spin $1/2$ primaries and a super-current. This is nothing but the third $\mathcal{N}=1$ super-Virasoro minimal model. and we therefore have only scalar primaries with $\Delta_{(\ell_{2},\ell_{3},\ell_{5})}=2h_{(\ell_{2},\ell_{3},\ell_{5})}$, i.e. a diagonal modular invariant. [^11] This can also be seen from the fact that at level $k=2$ as well as for odd levels $k=2n+1$ only the diagonal $A$-type modular invariant exists for $SU(2)$ WZW models Cappelli:1986hf . We list the first few operators in Table 2.
Table 2: Leading scaling dimensions in the candidate coset model.
@@
They turn out to roughly agree with the DMRG spectrum. Additionally, we can also analyze the $\mathbb{Z}_{2}$ quantum numbers of the operators by looking at the fusion rules of the coset primaries. Each member of the triplet satisfies its respective $SU(2)_{k}$ fusion rule, which is given by
$\ell_{k}\times\ell_{k}^{\prime}=\sum_{s_{k}=|\ell_{k}-\ell_{k}^{\prime}|}^{\textrm{min}(\ell_{k}+\ell_{k}^{\prime},\,k-\ell_{k}-\ell_{k}^{\prime})}s_{k}\,.$ (39)
The fusion rules are consistent with the diagonal $\mathbb{Z}_{2}$ symmetry, but don’t particularly illuminate the matching of operators. More importantly, considering the diagonal coset partition function, which has no accidental primary degeneracies and no additional multiplicities from the decomposition to Virasoro characters, it becomes clear that this coset does not have two-fold degeneracies for the light scalar operators and is therefore in tension with our DMRG spectrum.
There are, therefore, a few plausible scenarios:
  • The transition is weakly first order.
  • The CFT is rational but does not correspond to any known construction. This seems unlikely due to the relatively low value of the central charge.
  • The CFT is irrational (with potentially a positive twist gap).
We conjecture that the last scenario is realized. To give more definitive evidence for this claim, we can use the fact that rational CFTs with $c>1$ must have additional conserved currents other than the stress tensor. Therefore, accessing the spectrum of spinning operators with $J\neq 2$ and comparing it to the expectation of the existence of a single Virasoro vacuum module of conserved operators is a sharp criterion for irrationality. However, accessing the spectrum of these operators is challenging with vanilla DMRG techniques and requires some version of DMRG where one can preserve the momentum quantum number. We leave this analysis for future work.
2.2.1 Comments on the Landau-Ginzburg formulation
Before moving on to the three chain system, let us briefly comment on a possible Landau-Ginzburg realization of the same CFT. As is well known, the diagonal Virasoro minimal models $\mathcal{M}_{m,m+1}$ can also be realized as the IR fixed point of a non-compact scalar with a multi-critical potential Zamolodchikov:1987ti
which can be studied in the $3-\epsilon$ expansion. One could worry about negative directions in the potential, particularly for $N=2$, when taking, say $\phi_{1}\to-\infty,\phi_{2}\to\infty$ with $\phi_{1}/\phi_{2}$ fixed. Amusingly, our lattice deformation seems to know this. It is necessary to turn on the irrelevant operator $\epsilon^{\prime\prime}_{1}+\epsilon^{\prime\prime}_{2}$ to reach the critical point. This corresponds to the following Lagrangian, which is bounded from below:
$\mathcal{L}^{\prime}_{N-\rm{coupled}}=\int d^{2}x\sum_{i=1}^{N}\left(\frac{1}{2}(\partial\phi_{i})^{2}+\lambda_{1}:\phi_{i}^{6}:+\lambda_{3}:\phi_{i}^{8}:\right)+\lambda_{2}\sum_{i<j}^{N}:\phi_{i}^{3}\phi_{j}^{3}:\,.$ (42)
It would be interesting to see if this connection can be made precise and tested against results of the $\epsilon-$expansion Osborn:2017ucf ; Osborn:2020cnf .
3 3-coupled anyon chains
For the 3-chain model, we focus on the $K/J>0$ region to study the $N=3$ generalization of the CFT discussed in the previous section. In Fig.10, we give the Entanglement entropy measured at various values of $K/J$.
Figure 10: Entanglement Entropy as a function of the length of the segment. We have discarded the data points when the $l_{\rm conformal}$ is too small.
We can again set $K/J=\infty$ to measure the central charge, we get
$c=1.54\pm 0.01.$ (43)
We emphasize here that the error reported above is the standard error of the linear-fit parameter. The central charge of the CFT can be (slightly) smaller than the above value. We can also calculate the energy of excited states. To convert the energy spectrum into scaling dimensions, we again to calculate $E_{\infty}$ and the effective “speed of light” $v$ by studying how the ground state energy changes as a function of the lattice size. At $K/J=\infty$, the result is summarized in Fig. 11.
Figure 11: The average ground state energy as a function of the lattice size.
From this, we get the relation between energy and scaling dimension
$E\approx-1.825\leavevmode\nobreak\ L+\frac{13.72}{L}(\Delta-0.128).$ (44)
The scaling dimensions of the leading few CFT operators are summarized in Table 3.
$S_{3}$-irrep $Z_{2}$-charge $\Delta$
1singlet $+$ 0
2standard -- 0.049
1singlet -- 0.049
2standard $+$ 0.074
1singlet $+$ 0.074
Table 3: The scaling dimensions of the leading operators of the 3-chain CFT.
We now attempt to list candidate RCFTs as we did for the two-chain case. While no rigorous classification of RCFTs in increasing order of $c>1$ exists, [^12] An orthogonal classification scheme, where one instead increases the number of chiral primaries has produced interesting results Mathur:1988na ; Mukhi:2022bte , but is not sufficient for our purposes. all known RCFTs can be built as cosets of WZW models or obtained from Narain moduli spaces Yin:2017yyn . We will use two major classes of constructions to generate infinite families of RCFTs. These are cosets of the type
$\frac{g_{k}}{U(1)^{r_{g}}}\,,\qquad\frac{g_{k}\oplus g_{l}}{g_{k+l}}\,,$ (45)
where $r_{g}$ denotes the rank of $g$ [^13] This list is not claimed to be exhaustive in general, but we believe it is exhaustive for the two-chain case $c\leq 1.2$. For example, more general cosets of the type $G_{k}/H_{k}$ with $H$ a subgroup of $G$ such as $G=SU(N)$ and $H=SU(M)$ with $M<N$ also exist but have quickly growing central charges. There are also exceptional conformal embeddings (the so-called Maverick cosets Dunbar:1993hr ) which are believed to exist only for finitely values of the rank, but all known cases have central charges coinciding with the two main families discussed above.
Unfortunately, the sparseness of candidates found in the two chain case quickly subsides as we increase the central charge. In fact, we find several accumulation points from below on the central charge. We now list these infinite families of theories along with some finite number of exceptional cases. Again, we always keep in mind possible additional orbifolds of these theories.
  • Multiple accumulation families as $c\to 2$: $\mathcal{W}_{3}$ minimal models ($SU(3)_{1}\times SU(3)_{k}/SU(3)_{k+1}$) with $c=2(1-\frac{12}{(k+3)(k+4)})$; $SU(2)_{4}\times SU(2)_{k}/SU(2)_{k+4}$ cosets with $c=2(1-\frac{12}{(k+2)(k+6)})$; $\mathbb{Z}_{k}$ parafermions ($SU(2)_{k}/U(1)$) with $c=2(1-\frac{3}{k+2})$; Type II parafermions Dotsenko:2003ui ; Dotsenko:2003zc ($SO(3)_{2}\times SO(3)_{k}/SO(3)_{k+2}$) with $c=2(1-\frac{3}{(k+1)(k+3)})$; Infinite number of CFTs at $c=2$: Narain moduli space and orbifolds thereof.
  • $SU(3)_{2}\times SU(3)_{2}/SU(3)_{4}$ with $c=64/35\approx 1.8286$ but has same $c$ as previously mentioned coset. $SU(4)_{1}\times SU(4)_{k}/SU(4)_{k+1}$ with $k=1,2,3$ where once again the values of $c$ have appeared before. The same holds for $SU(5)_{1}\times SU(5)_{k}/SU(5)_{k+1}$ with $k=1,2$ and $SU(6)_{1}\times SU(6)_{1}/SU(6)_{2}$ but $SU(6)_{1}\times SU(6)_{2}/SU(6)_{3}$ with $c=25/12\approx 2.0833$ is new. All remaining $SU(N)_{1}\times SU(N)_{1}/SU(N)_{2}$ have $c<2$ but with same value as parafermions Bais:1987zk .
An obvious absence from the list are tensor product of minimal models with $c=2-\frac{6}{n(n+1)}-\frac{6}{m(m+1)}$ some of which can coincide with central charge values of previously listed cosets but most of which don’t (there are accumulation points for every $c=1+6/(m(m+1))$ and even accumulation points of accumulation points at $c=2$!). However, such tensor product theories have multiple stress tensors and are therefore forbidden on physical grounds. More generally, cosets of the form
$\frac{g_{k}\oplus g_{l}\oplus g_{m}}{g_{k+l+m}}\,,$ (46)
and its multi-copy generalizations, always have multiple stress-tensors and are again excluded.
The upshot from this analysis is that none of the rational theories listed above realize our non-invertible symmetry in any obvious way. It is therefore natural to conjecture once again that the $N=3$ model again flows to an irrational CFT.
4 Future directions
In this work, we have provided evidence for the existence of compact irrational CFTs in the phase diagram of $N=2,3$-coupled tricritical Ising models. The strongest evidence came in the $N=2$ case, where we used the knowledge of RCFTs with a similar central charge, along with their symmetries, to essentially rule out all the known candidate theories. Of course, there is no rigorous classification of RCFTs with $c>1$ and a more robust method is needed. As discussed previously, such RCFTs must have conserved currents other than the stress-tensor, and therefore developing a DMRG analysis at fixed momentum is key for excluding the existence of currents with low spin. [^14] More generally, understanding the space of consistent chiral algebras, and in particular how large the spin of the leading non-trivial current can be would be a very useful benchmark in this process. A particularly useful algorithm is the periodic uniform MPS method Zou:2017zce ; PhysRevB.83.125104 , which allows one to study states with periodic boundary conditions and fixed momentum. Furthermore, we argued that in the phase diagram of our model, a second, unstable CFT should be reachable by tuning an additional $\mathcal{C}$-invariant coupling. Studying this generalized lattice model would also allow us to tune away the leading irrelevant corrections to our stable fixed point, improving the numerical accuracy, and is therefore a worthy endeavor. This trick has proven to be effective in both Monte Carlo simulations PhysRevB.65.144520 ; PhysRevB.63.214503 and the recent fuzzy sphere regularization approach to study CFTs Zhu:2022gjc .
A closely related system to consider, are coupled 3-state Potts models, whose classical lattice version was studied in Dotsenko:1998gyp . While the standard local Hilbert space in which this system is formulated is 3-dimensional, which should further complicate our numerics, it is possible to flow to the same fixed point in a locally 2-dimensional Hilbert space Jacobsen:2024jel , by making use of the Temperley-Lieb algebra. Redoing our analysis in that setup seems promising, since other methods have been used in the past and are available for comparison, namely the $(Q-2)$-state expansion at fixed $N$, which makes it a perturbatively controlled fixed point, unlike the model we considered here.
Finally, we should mention other non-perturbative methods that are available to study our system, and that could complement our analysis. It will be interesting to explore the possibility of studying our coupled anyons Hamiltonian using quantum Monte Carlo simulations. This approach has the advantage that it would allow us to reach much larger system sizes, and might even allows us to consider the generalization of our model to $N\geq 4$ coupled chains. Orthogonally, we can also try to study the CFT in the IR directly, by making use of bootstrap methods. In two dimensions, two versions of the bootstrap could be useful: The sphere four-point bootstrap, [^15] See Kousvos:2024dlz for a bootstrap study of the coupled Potts CFT. where it is possible to impose the invertible global symmetry as well as some gaps motivated by our DMRG study; and the modular bootstrap, which studies the partition function on the torus, and where non-invertible symmetries can be implemented directly Lin:2023uvm .
Acknowledgements
We thank Connor Behan, Eduardo Castro, Clément Delcamp, Jesper Jacobsen, Igor Klebanov, Shota Komatsu, Sylvain Lacroix, João Lopes dos Santos, John McGreevy, Vítor Pereira, Francesco Russo, Slava Rychkov, Julian Sonner, Balt van Rees and Yifan Wang for useful discussions. AA further thanks Connor Behan for collaboration on related topics and Andreia Gonçalves for continued support. This work was granted access to the HPC resources of IDCS support unit from Ecole polytechnique. AA is funded by the European Union (ERC, FUNBOOTS, project number 101043588) and was further funded by the project 2024.00230.CERN at the University of Porto. JR is funded by the European Union (ERC “QFTinAdS”, project number 101087025). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them.
Appendix A RG instability of $SU(2)_{3}\times SU(2)_{3}/SU(2)_{6}$ coset CFT

### v3

@@ 4 Comparison to known RCFTs
  • $c=8/7\approx 1.14$ corresponding to the $\mathbb{Z}_{5}$ parafermion and its bosonized coset $SU(2)_{5}/U(1)$ Fateev and Zamolodchikov (1985); Dotsenko et al. (2003b).
@@
  • $c=81/70\approx 1.15$ corresponding to the third $\mathcal{N}=1$ super-Virasoro minimal model and its bosonized coset $SU(2)_{2}\times SU(2)_{3}/SU(2)_{5}$ Goddard et al. (1986).
  • $c=6/5=1.2$ corresponding to the second $\mathcal{W}_{3}$ minimal model Zamolodchikov (1985) ($SU(3)_{1}\times SU(3)_{2}/SU(3)_{3}$), as well as the coset $SU(3)_{2}/U(1)^{2}$ and the tensor product theory IsingTricritical $\textrm{Ising}\otimes\textrm${Tricritical Ising}.
Unfortunately, the sparseness of theories found for $c\leq 1.2$ quickly subsides as we increase the central charge. In fact, we find several accumulation points from below in $c$. Below, we list these infinite families of theories along with some finite number of exceptional cases. We should always keep in mind possible additional orbifolds of these theories and/or their non-diagonal modular invariants.
  • Multiple accumulation families as $c\to 2$: $\mathcal{W}_{3}$ minimal models ($SU(3)_{1}\times SU(3)_{k}/SU(3)_{k+1}$) with $c=2(1-\frac{12}{(k+3)(k+4)})$; $SU(2)_{4}\times SU(2)_{k}/SU(2)_{k+4}$ cosets with $c=2(1-\frac{12}{(k+2)(k+6)})$; $\mathbb{Z}_{k}$ parafermions ($SU(2)_{k}/U(1)$) with $c=2(1-\frac{3}{k+2})$; Type II parafermions Dotsenko et al. (2004, 2003a) ($SO(3)_{2}\times SO(3)_{k}/SO(3)_{k+2}$) with $c=2(1-\frac{3}{(k+1)(k+3)})$; Infinite number of CFTs at $c=2$: Narain moduli space and orbifolds thereof.
  • $SU(3)_{2}\times SU(3)_{2}/SU(3)_{4}$ with $c=64/35\approx 1.8286$ but has same $c$ as previously mentioned coset. $SU(4)_{1}\times SU(4)_{k}/SU(4)_{k+1}$ with $k=1,2,3$ where once again the values of $c$ have appeared before. The same holds for $SU(5)_{1}\times SU(5)_{k}/SU(5)_{k+1}$ with $k=1,2$ and $SU(6)_{1}\times SU(6)_{1}/SU(6)_{2}$ but $SU(6)_{1}\times SU(6)_{2}/SU(6)_{3}$ with $c=25/12\approx 2.0833$ is new. All remaining $SU(N)_{1}\times SU(N)_{1}/SU(N)_{2}$ have $c<2$ but with same value as parafermions Bais et al. (1988).
We will now argue that none of these models can realize a $((\rm{Fibonacci})^{3}\rtimes S_{3})$ symmetry in a straightforward manner (i.e. via Verlinde lines). In a coset model, the fusion rules are determined by the fusion rules of the primaries in the constituent WZW models. Using this, it is easy to convince one-self that the only way to obtain a line with a Fibonacci fusion rule is to already have such a line at the level of the individual factors. To further endow these lines with an $S_{N}$ permutation structure, we need to be able to swap the factors in the coset, but this is only possible for the two factors in the numerator, and only when they are equal. We therefore conclude that we can only straightforwardly realize $((\rm{Fibonacci})^{N}\rtimes S_{N})$ symmetry via cosets when $N=2$. We will further comment on this in Section 4.1 below.
An obvious absence from the list above are tensor product of minimal models with $c=2-\frac{6}{n(n+1)}-\frac{6}{m(m+1)}$ some of which can coincide with central charge values of previously listed cosets but most of which don’t (there are accumulation points for every $c=1+6/(m(m+1))$ and even accumulation points of accumulation points at c=2c=2! [^13] See Belin et al. (2025) for an interesting perspective on this point.). Such tensor product theories have multiple stress tensors. More generally, cosets of the form
$\frac{g_{k}\oplus g_{l}\oplus g_{m}}{g_{k+l+m}}\,$, (55)
and its multi-copy generalizations, where $((\rm{Fibonacci})^{N}\rtimes S_{N})$ topological lines could naturally be realized, always have $N$ stress-tensors. In our case however, if our CFT were a tensor product of three identical minimal models, the central charge of that theory would be below $c=0.7$ and above $c=0.5$, and such theories do not exist.
The upshot from this analysis is that none of the rational theories listed above realize our non-invertible symmetry in any obvious way. It is therefore reasonable to conjecture that the $N=3$ model flows to an irrational CFT.
4.1 CFTs with $(\rm{Fibonacci})^{2}\rtimes\mathbb{Z}_{2}$ symmetry
Our $N=2$ models preserve a $(\rm{Fibonacci})^{2}\rtimes\mathbb{Z}_{2}$ symmetry, and as we argued above, it is straightforward to realize such a symmetry in a coset CFT
$\frac{G_{k}\times G_{k}}{G_{2k}}\,$, (56)
if the WZW model $G_{k}$ already contains a Fibonacci Verlinde line $W$ associated to a primary in a representation $\rho_{k}$ and if $(\rho_{k},0,0)$ is allowed in the coset. We are aware of four WZW models where such $W$ lines are present
$G_{k}=SU(2)_{3}\,,\quad SU(3)_{2}\,,\quad(G_{2})_{1}\,,\quad(F_{4})_{1}\,$, (57)
respectively for the primaries in the representations
$\rho_{k}=(\mathbf{3})_{3}\,,\quad(\mathbf{8})_{2}\,,\quad(\mathbf{7})_{1}\,,\quad(\mathbf{26})_{1}\,$, (58)
where the boldface numbers denote the dimension of the representation. When we consider the coset (56) for the last two groups we get central charges corresponding to Virasoro minimal models with $c<1$. Instead, for $SU(2)_{3}$ we get the coset with $c=1.35$ identified with CFT2, which has also has an additional diagonal $\mathbb{Z}_{2}$ symmetry. Finally, for $SU(3)_{2}$ we get $c\approx 1.83$ as well as a diagonal $S_{3}$ symmetry, which agrees with the symmetries of our unknown CFT1. Unfortunately, the value of the central charge seems to be strictly outside our error bars for $c_{1}\approx 1.77$.
5 Future directions
In this work, we have conjectured the existence of compact irrational CFTs in the phase diagram of $N=3$ coupled tricritical Ising models by studying coupled anyonic chains with DMRG. Some indirect evidence for irrationality was given using the knowledge of RCFTs with a similar central charge, along with their symmetries, to essentially rule out all the known candidate theories. Of course, there is no rigorous classification of RCFTs with $c>1$ and a more robust method is needed. As discussed previously, such RCFTs must have conserved currents other than the stress-tensor, and therefore performing a DMRG analysis at fixed momentum is key for excluding the existence of currents with low spin. [^14] More generally, understanding the space of consistent chiral algebras, and in particular how large the spin of the leading non-trivial current can be would be a very useful benchmark in this process. A particularly useful algorithm is the periodic uniform MPS method Zou et al. (2018); Pirvu et al. (2011), which allows one to study states with periodic boundary conditions and fixed momentum.
In addition to $N$ copies of the tricritical Ising model coupled through the $\sigma^{\prime}$ operator (5), we have also studied the $N$ copies of the 3-state Potts models coupled through the $Z$ operator (7). In the large $N$ limit, such deformations can be viewed as double-trace deformations of the decoupled theory Giombi et al. (2018). Depending on whether the double-trace deformation is relevant or irrelevant, it leads to renormalization group flow towards a CFT in UV or IR. Classic examples are the flow between free scalar theory and the critical O(N) CFT, and the flow between free fermion theory and the critical Gross-Neveu model. We expect these corresponding CFTs of (5) and (7) to exist at least in the large $N$ limit. We have also shown here that the CFTs exist at $N=3$ for the tricritical Ising model and $N=2$ for the coupled Potts model. This leads us to conjecture that the CFT exists for any larger $N$, giving us two families of irrational CFT candidates.
A closely related system to consider, is the 3-state Potts models coupled by the energy operator, whose classical lattice version was studied in Dotsenko et al. (1999). While the standard local Hilbert space in which this system is formulated is 3-dimensional, which should further complicate our numerics, it is possible to flow to the same fixed point in a locally 2-dimensional Hilbert space Jacobsen and Wiese (2024), by making use of the Temperley-Lieb algebra. Redoing our analysis in that setup seems promising, since other methods have been used in the past and are available for comparison, namely the (Q2)(Q-2)-state expansion at fixed $N$, which makes it a perturbatively controlled fixed point. In fact, treating $Q$ as a continuous parameter allows us to study a family of fixed points containing both the examples considered in this paper and in Dotsenko et al. (1999), as well as potentially infinitely more examples corresponding to coupled minimal models of higher central Antunes et al. .
In this paper, we focus on a lattice model with $\mathcal{C}$-invariant coupling. Considering a more general lattice model with an additional $\mathcal{C}$-invariant coupling would allow us to tune away the leading irrelevant corrections to the fixed points, improving the numerical accuracy, and is therefore a worthy endeavor. This trick has proven to be effective in both Monte Carlo simulations Campostrini et al. (2002, 2001) and the recent fuzzy sphere regularization approach to study CFTs Zhu et al. (2023).
Finally, we should mention other non-perturbative methods that are available to study our system, and that could complement our analysis. It will be interesting to explore the possibility of studying our coupled anyons Hamiltonian using quantum Monte Carlo simulations. This approach has the advantage that it would allow us to reach much larger system sizes, and might even allows us to consider the generalization of our model to $N\geq 4$ coupled chains. Orthogonally, we can also try to study the CFT in the IR directly, by making use of bootstrap methods. In two dimensions, two versions of the bootstrap could be useful: The sphere four-point bootstrap, [^15] See Kousvos et al. (2024) for a bootstrap study of the coupled Potts CFT. where it is possible to impose the invertible global symmetry as well as some gaps motivated by our DMRG study; and the modular bootstrap, which studies the partition function on the torus, and where non-invertible symmetries can be implemented directly Lin and Shao (2023). [^16] See also Nakayama and Onagi (2025) for a first four-point bootstrap study using conditions motivated by non-invertible symmetry. We are currently actively investigating this direction.
Acknowledgements
We thank Connor Behan, Nathan Benjamin, Eduardo Castro, Luke Corcoran, Clément Delcamp, Paul Fendley, Jesper Jacobsen, Igor Klebanov, Shota Komatsu, Sylvain Lacroix, João Lopes dos Santos, John McGreevy, Vítor Pereira, Francesco Russo, Slava Rychkov, Hubert Saleur, Julian Sonner, Balt van Rees and Yifan Wang for useful discussions. AA further thanks Connor Behan for collaboration on related topics and Andreia Gonçalves for continued support. This work was granted access to the HPC resources of IDCS support unit from Ecole polytechnique. AA is funded by the European Union (ERC, FUNBOOTS, project number 101043588) and was further funded by the project 2024.00230.CERN at the University of Porto. JR is funded by the European Union (ERC “QFTinAdS”, project number 101087025). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them.
Appendix A Review of $SU(2)$ coset CFTs
In this appendix, we give a brief overview of the properties of the $SU(2)$ coset CFTs
$\frac{SU(2)_{k}\times SU(2)_{m}}{SU(2)_{k+m}}\,.$ (59)
The coset primaries are labeled by triplets of $SU(2)$ representations $(\ell_{k},\ell_{m},\ell_{k+m})$, where $\ell_{p}\in\{0,1/2,\dots,p/2\}$, subject to the identification $(\ell_{k},\ell_{m},\ell_{k+m})\equiv(k/2-\ell_{k},m/2-\ell_{m},(k+m)/2-\ell_{k+m})$. Furthermore, not all triplets correspond to allowed representations, and are further subject to the constraint
$\ell_{k}+\ell_{m}-\ell_{k+m}\in\mathbb{Z}\,.$ (60)
As is well known, the stress-tensor of the coset model is given by the difference of the numerator and denominator theories’ stress tensor (i.e. $T_{\textrm{coset}}=T_{k}+T_{m}-T_{k+m}$), and therefore the chiral dimensions of coset primaries satisfy
$h_{(\ell_{k},\ell_{m},\ell_{k+m})}=h_{\ell_{k}}+h_{\ell_{m}}-h_{\ell_{k+m}}+n\,,n\in\mathbb{N}\,$, (61)
where the addition of a positive integer might be needed since the primary does not always appear in the ground state of the denominator representation. Here, $h_{\ell_{k}}$ are the weights of primaries in the $SU(2)_{k}$ WZW model Wess and Zumino (1971); Witten (1983) and are given by
$h_{\ell_{k}}=\frac{\ell_{k}(\ell_{k}+1)}{k+2}\,.$ (62)
Given the finite list of chiral primaries, we can try to build different modular invariants for the coset theory. For example, if we observe sets of chiral dimensions which differ by an integer, a non-diagonal modular invariant might exist. A large set of modular invariants can be obtained by using the $ADE$-classified modular invariants of each constituent $SU(2)$ factor. Let us consider for example the coset
@@ Appendix A Review of $SU(2)$ coset CFTs
which has $c\approx 1.15$. Since for $k=2$ and $k$ odd, there is only an $A$-type modular invariant for $SU(2)_{k}$, we see that the coset does not admit non-diagonal modular invariants. Therefore, we can only have scalar primaries with $\Delta_{(\ell_{2},\ell_{3},\ell_{5})}=2h_{(\ell_{2},\ell_{3},\ell_{5})}$, i.e. a diagonal modular invariant. In fact, it is easy to see no integer differences between weights occur, [^17] Instead, one does find half-integer differences, and in particular a chiral dimension $3/2$. This is because the chiral theory admits a fermionic partition function with spin $1/2$ primaries and a super-current. This is nothing but the third $\mathcal{N}=1$ super-Virasoro minimal model. as can be seen by listing the first few operators as we did in Table 2.
Table 2: Leading scaling dimensions in the coset model (63).
@@
Finally, we describe how to fix the fusion rules. Each member of the triplet describing the primary satisfies its respective $SU(2)_{k}$ fusion rule, which is given by
$\ell_{k}\times\ell_{k}^{\prime}=\sum_{\ell_{m}=|\ell_{k}-\ell_{k}^{\prime}|}^{\textrm{min}(\ell_{k}+\ell_{k}^{\prime},\,k-\ell_{k}-\ell_{k}^{\prime})}\ell_{m}\,.$ (64)
As a representative example, we see that for $SU(2)_{3}$ the primary associated to the spin-1 representation of $SU(2)$ satisfies
$(1)_{3}\times(1)_{3}=(0)_{3}+(1)_{3}\,$, (65)
which is Fibonacci-like and will lead to actual Fibonacci topological lines in the case of the $SU(2)_{3}\times SU(2)_{3}/SU(2)_{6}$ coset explained below.
A.1 RG instability of $SU(2)_{3}\times SU(2)_{3}/SU(2)_{6}$ coset CFT